gbm1105:

Alina is able to swim at a rate of about 2.5 miles per hour. Talia is able to swim at a rate of 1.8 meters per second. Which statement comparing the two swimmers is accurate?

Mercury:

choices: A) Alina swims about 0.7 miles per hour faster than Talia. B) Talia swims about 0.7 miles per hour faster than Alina. C) Alina swims about 1.5 miles per hour faster than Talia. D) Talia swims about 1.5 miles per hour faster than Alina.

Mercury:

to compare the speeds, convert them both to the same units for example, here is the setup for converting Alina's speed in mph to meters per second, such that both speeds will be in meters per second \[\frac{ 2.5~miles }{ 1~hr }\frac{ 1609.34~meters }{ 1~mile }\frac{ 1~hr }{ 3600~seconds }\]